OpenStudy (anonymous):

PLEASE HELP!!!! 17. what are the solutions of the equation z^2-6z-27=0 A. 3, 9 B. -3, 9 C. 3, -9 D. -3, -9

OpenStudy (anonymous):

C

OpenStudy (anonymous):

a + b = -6 a * b = -27

OpenStudy (anonymous):

what would make the value of a and b in both of these equation true

OpenStudy (anonymous):

3-9 = -6 3 * -9 = -27